### v1

# Rethinking Few-Shot Fusion: Granular Ball Priors Enable General-Purpose Deep Image Fusion

Abstract
In image fusion tasks, due to the lack of real fused images as priors, most deep learning-based fusion methods obtain global weight features from original images in large-scale data pairs to generate images that approximate real fused images. However, unlike previous studies, this paper utilizes Granular Ball adaptation to extract features in the brightness space as priors for deep networks, enabling the fusion network to converge quickly and complete the fusion task. This leads to few-shot training for a general image fusion network, and based on this, we propose the GBFF fusion method. According to the information expression division of pixel pairs in the original fused image, we classify pixel pairs with significant performance as the positive domain and non-significant pixel pairs as the boundary domain. We perform split inference in the brightness space using Granular Ball adaptation to compute weights for pixels that express information to varying degrees, generating approximate supervision images that provide priors for the neural network in the structural brightness space. Additionally, the extracted global saliency features also adaptively provide priors for setting the loss function weights of each image in the network, guiding the network to converge quickly at both global and pixel levels alongside the supervised images, thereby enhancing the expressiveness of the fused images. Each modality only used 10 pairs of images as the training set, completing the fusion task with a limited number of iterations. Experiments validate the effectiveness of the algorithm and theory, and qualitative and quantitative comparisons with SOTA methods show that this approach is highly competitive in terms of fusion time and image expressiveness. Our code is now publicly available at:
Index Terms:
Granular Ball, Image fusion, Few-shot, Deep learning
I Introduction
Due to the limitations of sensors, images obtained from a single sensor often have their constraints and cannot effectively capture all the information within the field of view [1] The purpose of image fusion is to combine the image information obtained from different sensors, merging images of different modalities to produce a more comprehensive and information-rich image. For example, visible light images at night cannot capture thermal source information in dark areas, while images obtained from infrared imaging devices can significantly enhance the representation of thermal sources[2] Since the fused image combines the advantages of multiple source devices, it has wide applications in fields such as medical imaging, security, target recognition, and image segmentation[3].
In past research, significant achievements have been made in areas such as multi-focus image fusion, multi-exposure image fusion, infrared and visible light image fusion, and medical image fusion. These methods are mainly divided into traditional methods and deep learning methods. Traditional methods primarily decompose images through algorithms to extract features for fusion at different scales and spaces[4] However, as the requirements for use in real environments increase, the design of algorithms has become increasingly complex. Consequently, deep learning methods have gradually become mainstream. Among these, deep learning methods can be categorized into CNN-based methods, GAN-based methods, Transformer-based methods, and methods that combine traditional approaches. These methods have achieved remarkable results in this field. However, there is a lack of real images as references in different modality fusion tasks. These methods often rely on large-scale datasets for multiple iterations to produce images that closely resemble real fused images, which can be time-consuming. This presents significant challenges in real-world applications, such as long deployment times and extended update cycles.
To address the aforementioned challenges, this paper innovatively introduces coarse-grained spheres into the field of image fusion. Locally, it adaptively generates pixel-level masks through coarse-grained spheres, thereby producing a pseudo-supervised image as a prior for the output image of the neural network. At the global level, it utilizes saliency classification to adaptively generate weights that guide the neural network to converge quickly to the expected fused image, thus achieving few-shot training and rapid deployment. In previous research, coarse-grained spheres have already made significant achievements in data mining and machine learning, which gives them great potential for application in the field of image fusion. Since the adjacency relationship features in the original fused image are difficult to extract completely, and the brightness space features of the original images dominate the fusion process, we decompose the images in the brightness space. We generate coarse-grained spheres in the brightness space of the original images, gradually splitting and extracting pixel pairs with significant performance as the positive domain, while calculating the weights of non-significant pixel pairs within the coarse-grained spheres. Therefore, the spheres can locally and adaptively find the weights of different pixel points. At the global level, the spheres extract pixel pairs that exhibit significant performance for the fusion task, allowing them to perceive input images from different modalities. They adaptively adjust the final mask weights for different modalities to obtain supervised images suitable for various modality fusion tasks. Additionally, the proportion of significant pixel pairs at the global level will influence the establishment of the loss function, guiding the rapid convergence of the neural network along with the supervised images. In summary, our contributions are as follows:
  • To our knowledge, this is the first time that the Granular Balle theory has been applied to the field of general image fusion. An adaptive Granular Ball algorithm has been designed that can not only compute pixel weights locally but also globally perceive the ratio of modalities to salient pixels.
  • A universal image fusion model has been successfully implemented for few-shot training, introducing a paradigm for few-shot training in image fusion. By introducing adaptive Granular Ball to provide priors, the model quickly and effectively achieved universal image fusion tasks using only 10 pairs of training data for each modality.
  • A neural network structure and loss function were designed based on the supervised images and significant feature point weights generated from Granular Ball. By utilizing the global confidence from the Granular Ball generation, a clever one-to-one weight setting was implemented for each pair of images during the training phase, which encourages the neural network to converge quickly, allowing the image fusion task to be completed in less than 100 iterations.
  • Compared to previous methods, the proposed method can complete image fusion tasks more quickly and efficiently. The effectiveness and rationality of the algorithm have been validated through experiments, and it has been compared with SOTA methods on public datasets, achieving competitive performance.
The structure of the following sections is as follows: In Section 2, we will briefly introduce image fusion methods and granular computing. In Section 3, we will provide a detailed description of the structure and algorithm implementation. In Section 4, we will conduct experiments for validation, compare with SOTA methods, and provide theoretical verification. Section 5 will summarize the entire text.
II Related Work
In this section, we will introduce the methods and developments of image fusion, as well as the theory and applications of granular computing.
II-A Fusion method
Image fusion requires methods that can extract features from different modality images and highlight the effective information from each modality after fusion, thereby fully representing the real scene [5]. In past research, researchers primarily used methods such as multi-scale transformations, subspace analysis, and sparse representation. As the application scenarios became increasingly complex, some researchers began to combine these methods to enhance information extraction capabilities[6]. Early researchers used Laplacian pyramids (LP) to decompose images across different frequency bands, and later utilized dual-tree complex wavelet transform (DWCWT) to extract effective information through different filtering for image fusion. However, wavelets have poor edge extraction capabilities, leading researchers to propose the curvelet transform method, which effectively captures edge information. To enhance the representation of structural information in fusion, guided filtering (GFF) [7] was also applied to image fusion. Additionally, Liu et al. combined sparse matrix representation with traditional transformation methods to achieve good results in extracting effective information from fused images[8].
However, in real-world environments, as the requirements for usage increase, image processing algorithms often exhibit poor generalizability, leading to increasingly complex designs and structures for these algorithms, which poses challenges for their application in real-world settings[6].With the widespread application of deep learning, Li et al. were the first to introduce deep learning into the field of image fusion[9], making it the mainstream approach in subsequent research. Networks with encoding and decoding structures are considered to have stronger detail extraction capabilities. The DeepFuse method combines encoding and decoding structures with the fusion of infrared and visible light images[10], demonstrating good performance. Building on this, there are methods that use pre-trained networks for feature extraction, leading to the development of the IFCNN method[11], which is suitable for general fusion. Utilizing pre-trained networks to extract features can yield a weight for image saliency, and later, U2fusion leveraged this characteristic to design a loss function that implicitly guides the fusion of images from different modalities during the iterative process[12]. Unlike CNN structures, Generative Adversarial Networks (GANs) use a discriminator network to assess the accuracy of each pixel generated by the generator network. The FusionGAN method was the first to apply this to the fusion of infrared and visible light images[13]. Building on this method, Ma et al. proposed the DDcGan method by adding more discriminators to enhance accuracy[14], achieving satisfactory results. The Transformer, known for its unique long-range dependencies in text processing[15], was also introduced into the field of image processing by Ma et al., who decomposed images into blocks to establish connections for image fusion, resulting in the proposal of SwinFusion[16]. Later research found that neural networks can also learn the logic of traditional methods, and combining traditional methods with deep learning approaches can yield higher quality results in image fusion. For instance, MMIF-INET successfully achieved medical image fusion by combining wavelet decomposition with reversible neural networks[17]. MetaFusion utilized meta-features for interactive information features to achieve fusion of infrared and visible light images[18], while ZMFF employed deep priors to generate depth masks for sample-free multi-focus image fusion[19][20]. Although these deep learning methods have achieved good results, practical applications, such as those based on CNNs and Transformers, still require large datasets to train models and extract parameters from original images[21][22]. Moreover, sample-free fusion methods like ZMFF can consume a significant amount of time to produce a fused image. Deep learning methods that rely on traditional methods as priors struggle to find an adaptive fusion scale, presenting challenges for rapid deployment and low-sample image fusion.
II-B Granular-ball computing
Rough sets are used to describe the uncertainty of problems, while ball granules simplify finer data items through granulation, described by radius and centroid in a Granular Ball model in space[23][24]. Later, scholars combined these concepts and proposed the GBNRS method to handle continuous data using granulation and approximation[25]. Based on this, they used Granular Ball clusters to describe the relationships in clustering, significantly reducing spatial complexity and proposing the Ball k-means method in the direction of clustering[26]. Ball granule computation has shown great potential in data processing and handling uncertainty relationships. However, in the field of image processing, due to the lack of a clear model establishment, previous ball granule methods have struggled to describe the relationships between images.
To address the aforementioned challenges, we will integrate images into a brightness subspace, extracting pixel pairs from the fused image that can directly exhibit significant performance during the fusion process. The characteristics of these pixel pairs are often rough and difficult to extract directly in the absence of prior knowledge. Therefore, we will categorize all pixel pairs with significant contributions during the fusion process into a positive domain, while those whose contributions are difficult to determine due to a lack of prior knowledge will be categorized into a boundary domain. All pixel pairs in the fused image are treated as a coarse-grained sphere, and in the brightness space, we continuously extract non-significant points of pixel pairs through finer-grained spheres. This is because non-significant pixel pairs often have similar information in the brightness space, while significant pixel pairs are extracted when the grain sphere reaches its limit and splits. Ultimately, we obtain pixel pairs described by the positive and boundary domains, where significant pixel pairs typically describe the spatial structural information of the image. The proportion of the positive domain will determine its confidence as a prior in the neural network, which will make decisions regarding the boundary domain pixel pairs, thereby determining all pixel pairs for image fusion. We will introduce the proposed methods and theories in detail in the following chapters.
Figure 1: The overall framework of GBFF, where the dashed parts in the diagram are algorithms and modules used only during the training phase, and the solid parts are those needed in both the testing and training phases.
III Proposed Methods
In this section, we will introduce the overall framework process according to the training phase and the testing phase, and then provide a detailed explanation of the principles of the Granular Ball swarm algorithm, the network structure, and the configuration of the loss function.
III-A Framework
As shown in the Fig.1, the structural diagram of our framework is fully presented, including the settings of each module in the network. It is important to note that the dashed parts of the structural diagram are modules used only during the training phase, while the solid parts are used in both the testing and training phases. Here, SP refers to significant pixel pairs, and NSP refers to non-significant pixel pairs. We have two input paths, A and B, which can be interchanged freely. All input images will be converted to the YCbCr color space. In the network structure section, the numbers on the input paths and output paths represent the number of feature map channels, and all convolution modules in the network use 3x3 convolution kernels. In this paper, the network emphasizes quickly fitting its own detail information, thus multiple Cat operations are used to concatenate feature maps to enhance the network’s ability to extract its own information.
III-B Granular Ball generation
The brightness space of the original fused image carries the majority of the information, where the brightness of the pixels plays a role in information expression. During the fusion process, the pixels of the two fused images appear in pairs, and these pixel pairs contain the information of the true fused image. A desired fused image is obtained by expressing its pixel information according to certain weights. It is clearly unreasonable to simply use a single metric to classify the significance of pixel pairs based on brightness information within a single scale. Decomposing a scale into multiple local scales can also lead to insufficient priors, making it difficult to infer whether pixel pairs play a significant role in the fusion. Therefore, we propose a method that differs from traditional Granular Ball generation. We gradually expand our perception of global information through the adaptive sliding of the Granular Ball boundaries, using the brightness information features of non-significant points as priors. In this process, we gradually extract pixel pairs with similar information, while pixel pairs that exhibit significance will be delayed until the Granular Ball captures enough information and reaches the boundary that requires splitting before extraction.
Definition 1. We define a granular ball $\mathcal{G}(\mu,r)$ in the brightness space as:
$\mathcal{G}(\mu,r)$ =$=\left\{x\in\mathbb{R}\,\middle|\,|x-\mu|\leq r\right\},$ $\mathcal{G}(\mu,r)\subseteq[0,255]$ (1)
Given a pair of grayscale images $A$ and $B$, a pixel at position $(x,y)$ falls within the granular ball if:
$\begin{array}[]{c}(x,y)\in\mathcal{P}(\mu,r)\\[5.0pt] \text{s.t.}\quad A_{x,y}\in\mathcal{G}(\mu,r)\lor B_{x,y}\in\mathcal{G}(\mu,r)\end{array}$ (2)
III-C Significant pixel extraction
We hope to estimate pixel pairs with significant information expression capability in the original image through Granular Ball, which only have a local perception domain. To achieve this, for two input images A and B, we first define pixel pairs at the local level.
Definition 2. Each pixel pair $(A_{x,y},B_{x,y})$ from the source images is categorized into :
NSP: Both pixels fall within the granular ball.
$\begin{array}[]{c}(A_{x,y},B_{x,y})\in\mathcal{NSP}\\[5.0pt] \text{s.t.}\quad A_{x,y}\in\mathcal{G}(\mu,r)\land B_{x,y}\in\mathcal{G}(\mu,r)\end{array}$ (3)
SP: Only one of the pixels lies within the granular ball.
$\begin{array}[]{c}(A_{x,y},B_{x,y})\in\mathcal{SP}\\[5.0pt] \text{s.t.}\quad A_{x,y}\in\mathcal{G}(\mu,r)\land B_{x,y}\notin\mathcal{G}(\mu,r)\\ \quad\lor A_{x,y}\notin\mathcal{G}(\mu,r)\land B_{x,y}\in\mathcal{G}(\mu,r)\end{array}$ (4)
Definition 3. Given a granular ball $\mathcal{G}(\mu,r)$, we define the purity of the granular ball as the proportion of non-salient pairs (NSP) within it:
$\text{Purity}(\mathcal{G})=\frac{|\mathcal{NSP}|}{|\mathcal{NSP}|+|\mathcal{SP}|}$ (5)
where, $|\mathcal{NSP}|$ and $|\mathcal{SP}|$ denote the number of pixel pairs classified as non-salient and salient, respectively, under the current granular ball $\mathcal{G}(\mu,r)$. Due to the initially generated Granular Ball having only a local receptive field, the priors in the Granular Ball balls are insufficient to infer pixel pairs that exhibit significant information. Therefore, we prioritize processing non-significant pixel pairs and delay the processing of pixel pairs in the positive domain. This occurs when the Granular Ball reach a certain state. Here, we define the purity of local Granular Balls, as well as the positive domain and boundary domain at the global level.
Definition 4.The pixel pairs in images are interpreted as follows:
  • The set of salient pairs (SP) corresponds to the positive region:
    $\mathcal{POS}(A,B)=\left\{(x,y)\,\middle|\,(A_{x,y},B_{x,y})\in\mathcal{SP}\right\}$ (6)
  • The set of non-salient pairs (NSP) corresponds to the boundary region:
    $\mathcal{BND}(A,B)=\left\{(x,y)\,\middle|\,(A_{x,y},B_{x,y})\in\mathcal{NSP}\right\}$ (7)
III-D Local weight calculation and global perception
As shown in the Fig.2, non-significant points are gradually extracted at the local level through coarse-grained spheres, and pixel pairs belonging to the boundary domain and positive domain are inferred at the global level. However, non-significant pixels that fall within the boundary domain often lack prior information to infer their weights during fusion. Therefore, we use a greedy strategy to roughly estimate and retain as much information as possible. The algorithm for the coarse-grained sphere gradually diffusing from local to global is shown in Algorithm 1.
    Algorithm 1:Granular Ball Significant Extraction
0:  Input images $A,B$
0:  Granular ball radius $r$, initial mask $X=0$, initial ranges $E_{1},E_{2}$
0:  Final fusion mask $X$, $\mathcal{SP}$, $\mathcal{NSP}$
1:  /* Function AssignSP$(E_{\text{low}},E_{\text{high}})$/
2:  for all $(x,y)\in\mathcal{SP}$ and $X_{x,y}=0$ do
3:     if $A_{x,y},B_{x,y}\in[E_{\text{low}},E_{\text{high}}]$ then
4:        if $A_{x,y}>B_{x,y}$ then
5:           $X_{x,y}=0.95$
6:        else
7:           $X_{x,y}=0.05$
8:        end if
9:     end if
10:  end for
11:  /* Function AssignNSP$(E_{\text{low}},E_{\text{high}})$/
12:  for all $(x,y)\in\mathcal{NSP}$ and $X_{x,y}=0$ do
13:     if $A_{x,y},B_{x,y}\in[E_{\text{low}},E_{\text{high}}]$ then
14:        $D_{1}=|\max(A_{x,y},B_{x,y})-E_{\text{high}}|$
15:        $D_{2}=|\max(A_{x,y},B_{x,y})-E_{\text{low}}|$
16:        $\omega=\max(D_{1},D_{2})/((E_{\text{high}}-E_{\text{low}})+10^{-6})$
17:        $\omega=\min(\omega,0.95)$
18:        if $A_{x,y}\geq B_{x,y}$ then
19:           $X_{x,y}=\max(\omega,0.05)$
20:        else
21:           $X_{x,y}=1-\max(\omega,0.05)$
22:        end if
23:     end if
24:  end for
25:  /* Main Procedure */
26:  Initialize $\mathcal{SP}=\varnothing$, $\mathcal{NSP}=\varnothing$
27:  $E_{1}=255$, $E_{2}=E_{1}-2r$
28:  while there exist unassigned pixels in $X$ do
29:     if $(E_{1}-E_{2})>\beta\cdot r$ then
30:        $\text{step}=\lfloor(E_{1}-E_{2})/2\rfloor$
31:        $\text{rangeLow}=E_{1}-\text{step}$,  $\text{rangeHigh}=E_{1}$
[…]
33:        $E_{1}=\text{rangeLow}$
34:     end if
35:     while $E_{1}>E_{2}$ and no unprocessed pixel in $[E_{2},E_{1}]$ do
36:        $E_{1}=E_{1}-1$
37:        if $E_{1}\leq E_{2}$ then
38:           $E_{2}=\max(E_{2}-2r,0)$
39:           $E_{1}=E_{2}$
40:        end if
41:     end while
42:     Call AssignSP$(E_{2},E_{1})$
43:     Call AssignNSP$(E_{2},E_{1})$
44:     $E_{2}=\max(E_{2}-2r,0)$
45:     $\mathcal{SP}^{\prime},\mathcal{NSP}^{\prime}=\{\text{newly unprocessed pixels in }[E_{2},E_{1}]\}$
46:     $\text{Purity}=\dfrac{|\mathcal{NSP}^{\prime}|}{|\mathcal{SP}^{\prime}|+|\mathcal{NSP}^{\prime}|}$
47:     if $\text{Purity}=1$ or $E_{2}<\delta$ then
48:        break
49:     end if
50:  end while
51:  return  $X,\mathcal{SP},\mathcal{NSP}$
Where the NSP calculates weights within a local Granular Ball. The pixel points extracted from the NSP within the Granular Ball adaptively obtain positively correlated weights based on their own information content. For SP, inference is only performed after the termination signal is triggered, and a value of 0.95 is used to prevent information loss. The iteration of the Granular Ball uses boundary inference, gradually iterating the boundary range by utilizing the brightest and darkest pixel points within, allowing the Granular Ball to roll in the brightness space. During the boundary inference process, some SPs are also extracted. In the final stage, as shown in Algorithm 2, all pixel pairs are divided into positive and boundary domains. The ratio of pixel pairs in the positive domain is used to perceive the modality type of the image. The positive domain pixel pairs in multi-exposure images will be significantly higher than those of other modalities, and by directly correcting the weights of the pixel pairs (SP) in the positive domain, the pixel information in overexposed areas can be suppressed. It should be noted that for a pair of input images, the SP ratio is numerically equal to the POS ratio, but here SP and NSP emphasizes locality, while POS and BND emphasizes globality.
Figure 2: It describes the splitting of particles and the process of extracting significant pixels during the splitting process. After the Granular Ball split, they will continue to slide until a termination condition is met or they continue to split.
   Algorithm 2: Modal Perception
0:  Final mask $X$, Classification $NSP$, $SP$,Input images $A,B$
0:  Supervise image Y , Output $r_{\text{$BND$}}$,$r_{\text{$POS$}}$
1:  Count total number of pixels $N$ in $X$
2:  Compute $POS$ ratio: $r_{\text{$POS$}}=\frac{N_{\text{$SP$}}}{N}$
3:  $r_{\text{$BND$}}$=$1-r_{\text{$POS$}}$
4:  if $r_{\text{POS}}>0.95$ then
5:     Identify all $(x,y)$ from $POS$
6:     for all such pixels $(x,y)$ do
7:        $X_{x,y}=0.5$
8:        $r_{\text{$BND$}},r_{\text{$POS$}}$ = $0.5,0.5$
9:     end for
10:  end if
11:  $Y=X*A+(1-X)*B$
12:  return  $Y$, $r_{\text{$BND$}}$,$r_{\text{$POS$}}$
III-E Multi-source input fusion strategy
The neural network trained by our method has the ability to respond to salient information and edges, as shown in Fig.3. By sequentially inputting multi-source images into the fully trained neural network, the network can extract salient information, thereby generating an image fused with salient information at each fusion step, ultimately resulting in a fused image that contains information from multiple source images.
Figure 3: Mutiple inputs fusion process
III-F Loss Function
The fused image processed through Granular Ball treatment can obtain pixel-level masks. By using these masks for image fusion, we can obtain supervised images that provide structural priors. At the same time, SP and NSP are used to calculate $r_{POS}$, which is used to evaluate the information confidence of the supervised images. Therefore, Granular Ball saliency extraction provides supervised images, The Granular Ball provides supervised images, $r_{POS}$, and $r_{BND}$ during the training phase of the neural network, and the loss function of the neural network will be set around them. We describe the real fused image information as: (a) the spatial structural information of the supervised images, (b) the boundary information determined by $r_{POS}$ and the boundary information of the original image determined by $r_{BND}$, and (c) the high-frequency information of the original fused image. Therefore, we divide the loss function into three parts:
Figure 4: To facilitate the observation of the significance extraction of the granules in the image weight calculation, the image is divided into small blocks of 128*128. The gray areas in (e) Weight Mask represent the parts that have been adjusted by the weights, while (d) Supervise image is the prior supervised image obtained through the weight mask.
a) Structural Similarity Loss:
$\mathcal{L}_{\text{ssim}}=1-\text{SSIM}(O,S)$ (8)
$\text{SSIM}(O,S)=\frac{(2\mu_{O}\mu_{S}+C_{1})(2\sigma_{OS}+C_{2})}{(\mu_{O}^{2}+\mu_{S}^{2}+C_{1})(\sigma_{O}^{2}+\sigma_{S}^{2}+C_{2})}$ (9)
where $\mu_{O}$, $\mu_{S}$ are local means, $\sigma_{O}^{2}$, $\sigma_{S}^{2}$ are variances, and $\sigma_{OS}$ is the covariance. Constants $C_{1}$ and $C_{2}$ stabilize the division.
b) Sobel Gradient Loss:
To guide the fusion process with edge-aware gradient information, we use the Sobel operator to extract gradients:
$G_{x}=\begin{bmatrix}-1&0&+1\\ -2&0&+2\\ -1&0&+1\end{bmatrix},\quad G_{y}=\begin{bmatrix}-1&-2&-1\\ 0&0&0\\ +1&+2&+1\end{bmatrix}$ (10)
The gradient magnitude is computed as:
$\text{Sobel}(I)=\sqrt{G_{x}^{2}+G_{y}^{2}}$ (11)
Due to the proportion of significant pixel pairs extracted by the Granular Ball being POS, we define the boundary loss function as:
$G_{\text{target}}=r_{\text{$POS$}}\cdot\text{Sobel}(S)+r_{\text{$BND$}}\cdot\left(\text{Sobel}(X)+\text{Sobel}(Y)\right)$ (12)
where the $X$ and $Y$ are original fused images.The Sobel loss minimizes the L1 difference between the output gradient and this guided target.The second part of loss function is:
$\mathcal{L}_{\text{Sobel}}=\|\text{Sobel}(O)-\text{Sobel}(G_{\text{target}})|_{1}$ (13)
c) Laplacian Loss:
The Laplacian filter is applied to all input and target images to extract high-frequency energy information:
$\Delta=\begin{bmatrix}0&1&0\\ 1&-4&1\\ 0&1&0\end{bmatrix}$ (14)
Therefore, the third part of the loss function is:
$\mathcal{L}_{\text{Lap}}=\|\Delta T-\Delta X\|_{1}+\|\Delta T-\Delta Y\|_{1}\$ (15)
Finally, we obtain the loss function for the entire network:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{ssim}}+\mathcal{L}_{\text{Lap}}+\mathcal{L}_{\text{Sobel}}\$ (16)
IV Experimental Results And Analysis
IV-A Experimental Setup
For the task of fusing infrared and visible light images(VIF), we selected 5 pairs of images from the MSRS and TNO datasets as the training set. For medical image fusion(MIF), we selected 10 pairs of images from the Harvard (PET-MRI) dataset as the training set. For multi-exposure image fusion(MEF), we selected 10 pairs of images from the MEFB dataset as the training set. Finally, for the multi-focus image fusion task(MFF), we selected 10 pairs of images from the Lytro dataset as the training set. In the subsequent steps, we will use the images from the datasets that were not used in training as the test set. To validate generalizability, we will add the RoadScene dataset as a test set for the infrared and visible light image fusion task, and the MFI-WHU dataset as a test set for multi-focus fusion.
During the training phase, to enhance the robustness of the neural network, all input images will be cropped to 128x128, with a stride set to 64 and a batch size set to 16. We will also perform flipping and shuffling operations on the images in the batch to increase the sample size. Due to the higher resolution of multi-exposure images, they can be divided into more patches, so the epoch for multi-exposure images is set to 5, while the epoch for other modalities is set to 100. All experiments will be conducted on an RTX 4070 TIS platform, with a CPU of 14600KF.
TABLE I: Quantitative comparison of fusion performance across MSRS, RoadScene, and TNO datasets using five metrics.
Method MSRS RoadScene TNO
MI SD PSNR VIF Qab MI SD PSNR VIF Qab MI SD PSNR VIF Qab
BBD 2.210 28.435 19.745 0.342 0.380 2.681 44.405 14.508 0.254 0.370 2.497 44.377 15.475 0.224 0.265
DE 3.092 34.908 22.363 0.489 0.505 3.110 35.274 16.975 0.228 0.372 2.746 30.508 14.581 0.238 0.348
IFCNN 2.836 35.331 22.900 0.502 0.583 3.000 38.350 16.987 0.252 0.407 2.505 33.188 15.561 0.326 0.471
SD 0.570 27.364 12.118 0.019 0.026 0.988 36.904 12.451 0.026 0.034 2.661 26.462 11.163 0.023 0.046
SwIN 2.262 26.058 17.311 0.303 0.284 3.190 43.828 15.837 0.257 0.374 2.492 39.243 14.908 0.268 0.469
U2 2.262 26.058 17.311 0.303 0.284 3.025 39.830 14.945 0.195 0.390 2.110 28.594 15.665 0.281 0.337
MetaFusion 1.799 39.189 21.481 0.530 0.488 2.240 51.148 15.539 0.274 0.347 1.844 40.321 16.513 0.423 0.352
OURS 3.306 38.499 22.388 0.533 0.591 3.351 46.254 17.333 0.324 0.425 2.892 41.378 15.566 0.354 0.475
IV-B Multimodal image fusion
In this section, we will compare our method with the following state-of-the-art (SOTA) methods for general image fusion: (c) DDBFusion[27], (d) DEFusion[28], (e) IFCNN, (f) SDNET[9], (g) SwinFusion, and (h) U2Fusion. To demonstrate the effectiveness of our method, we will introduce a dedicated SOTA method for each type of modality image fusion, specifically: (i) MetaFusion for infrared and visible light fusion, (j) MMIF-INET for MRI and PET image fusion, (k) ZMFF for multi-focus image fusion, and (l) SAMT-MEF[29] for multi-exposure image fusion. In the quantitative analysis section, we will use the following evaluation metrics: MI (to assess the amount of information contained in the fused image), SD (to measure the contrast of the fused image), PSNR (to evaluate the reconstruction quality of the fused image), VIF (to assess the visual fidelity of the fused image), Qab (to evaluate the structural preservation of the fused image), EN (to assess the information entropy of the fused image), SF (to evaluate the spatial frequency of the fused image), and AG (to assess the clarity of the fused image). The higher the values of these evaluation metrics, the better the quality of the fused image. For ease of observation, we will use bold black to mark the optimal indicators and underline the sub-optimal indicators.
IV-B1 Infrared and visible image fusion
As shown in Fig.5, compared to the methods being compared, our method demonstrates a stronger advantage in the detail texture of visible light images and infrared features, resulting in fused images with less noise and more distinct boundaries. For quantitative analysis, we used MI , SD , PSNR, VIF, and Qab. The comparison results with various methods are shown in Table I. Our method shows a strong advantage in the amount of information, fidelity, and structural retention of the fused image, and it also achieved second place in other evaluation metrics. Overall, the fused images produced by our method demonstrate a more competitive performance, capable of completing fusion tasks with fewer samples, showcasing good generalization ability.
Figure 5: Select a typical image fusion result from TNO, RoadScene, and MSRS datasets for qualitative comparison.
IV-B2 PET and MRI image fusion
The comparison results are shown in Fig.6. The fused images obtained from GBFF demonstrate a stronger advantage in texture details and edge performance, preserving the features of the PET images while also showcasing the texture characteristics of the MRI images. As shown in Table II, the GBFF method exhibits optimal results in the quantitative analysis of EN , MI, and PSNR values, indicating an advantage in the information content and contrast of the fused images. Additionally, the suboptimal SD and Qab values also indicate that the fused images obtained by the GBFF method are correlated with the original fused images.
TABLE II: The metric scores obtained from evaluating each method on the Harvard (PET-MRI) dataset.
MIF Metrics
Method EN MI SD PSNR Qab
BBD 5.209 3.018 78.647 16.691 0.475
DE 5.183 2.922 77.806 19.374 0.575
IFCNN 5.041 2.890 82.178 19.933 0.629
SD 5.491 2.835 67.487 16.079 0.541
SwIN 5.233 2.548 83.406 18.956 0.600
u2 5.555 3.040 70.196 15.722 0.336
mmif 5.024 3.031 91.864 19.305 0.672
OURS 5.594 3.042 85.484 20.104 0.632
Figure 6: Select the fusion results of 3 typical images from the Harvard (PET-MRI) dataset for qualitative comparison.
IV-B3 Multi-focus image fusion
As shown in Fig.7, the GBFF method benefits from the significant extraction of granules, resulting in fused images with higher boundaries and contrast compared to other methods, leading to clearer representations. As indicated in Tables III and IV, quantitative analyses on two datasets show that the GBFF method achieves optimal values in EN, SD, and Qab for fused images, indicating that the fused images maintain high contrast and information content while preserving the original fusion image information. Additionally, in terms of SF and AG , it achieved the best and second-best results on the two datasets, respectively. Overall, the GBFF method demonstrates stronger advantages and competitiveness compared to other methods.
TABLE III: The metric scores obtained from evaluating each method on the Lytro dataset.
Lytro Metrics
Method EN SD SF AG Qab
BBD 7.535 57.153 0.023 7.556 0.189
DE 7.463 54.356 0.022 6.724 0.198
IFCNN 7.523 57.096 0.036 11.276 0.235
SD 7.482 55.270 0.033 10.369 0.244
SwIN 7.524 56.712 0.032 9.842 0.243
u2 7.460 58.390 0.023 7.073 0.207
zmff 7.524 56.954 0.035 11.003 0.059
OURS 7.611 61.989 0.037 11.526 0.245
TABLE IV: The metric scores obtained from evaluating each method on the MFI-WHU dataset.
MFI-WHU Metrics
Method EN SD SF AG Qab
BBD 7.312 52.557 0.033 10.464 0.035
DE 7.223 49.534 0.026 8.320 0.033
IFCNN 7.326 52.714 0.050 14.896 0.037
SD 7.288 51.255 0.045 14.744 0.038
SwIN 7.289 51.589 0.040 12.803 0.037
u2 7.226 53.326 0.026 8.230 0.032
zmff 7.295 51.994 0.045 14.780 0.035
OURS 7.419 58.100 0.046 14.792 0.039
Figure 7: Select the fusion results of 3 typical images from the MFI-WHU and Lytro datasets for qualitative comparison.
IV-B4 Multi-exposure image fusion
Compared to other fusion methods, as shown in Fig.8, the GBFF method for multi-exposure image fusion focuses more on the representation of texture information, avoiding texture distortion like that caused by the IFCNN method and color distortion like that caused by the SAMT-MEF method. Additionally, in high-exposure environments, the GBFF method handles light sources more smoothly, preserving details without causing halo noise. Quantitative analysis is shown in Table V, where the GBFF method achieves optimal results in EN and Qab, indicating that the fused images obtained by the GBFF method can capture more information during fusion. The suboptimal results in MI, SD, and SF suggest that the fused images have a high amount of information and contrast.
TABLE V: The metric scores obtained from evaluating each method on the MEFB dataset.
MEFB Metrics
Method EN MI SD SF Qab
BBD 6.730 0.722 58.406 0.027 0.028
DE 6.463 0.592 38.457 0.017 0.023
IFCNN 6.888 0.465 46.404 0.041 0.027
SD 6.623 0.668 40.202 0.023 0.025
SwIN 6.212 0.703 33.387 0.012 0.017
u2 6.739 0.744 39.674 0.013 0.022
SAMT 6.849 0.792 43.505 0.024 0.038
OURS 6.947 0.678 46.655 0.028 0.031
Figure 8: Select the fusion results of 3 typical images from the MEF dataset for qualitative comparison.
IV-C Efficiency Analysis
In this section, we will compare the efficiency of all methods using images of size 224*224 as a benchmark, and we will use the network’s training parameters (params), floating-point operations (Flops), and the time consumed to fuse a pair of images as metrics. As shown in Table X, due to the fewer residual layers in the GBFF network and its simplicity, although it did not achieve the best results in terms of params, it achieved the optimal value in Flops and also performed second best in fusion time. We compared the training time using the training method recommended by the authors and recorded the time consumed for model training, as shown in Table VI, which details the time consumed by GBFF for training across different modalities. Compared to U2Fusion and IFCNN, GBFF has a significant advantage in training time, allowing for faster deployment and usage.
TABLE VI: Comparison of parameter size, computational cost, and inference time among different image fusion methods.
Method Params (M) FLOPs (G) Time (s)
BBD 3.673 184.931 0.953
DE 7.874 71.551 0.426
IFCNN 0.083 39.932 0.024
SD 0.044 25.638 0.056
SWIN 0.973 292.537 1.467
U2 0.659 405.174 0.143
SAMT 1.176 38.450 0.174
MMIF 0.749 3.020 0.110
ZMFF 4.670 230.380 242.790
META 9.350 24.630 0.220
OURS 0.261 2.211 0.022
V Theoretical analysis and ablation experiments
V-A Analysis of Granular Ball Significant Extraction Algorithm
In this section, we will analyze the impact of the radius of the Granular Ball on the extraction of significant pixels. Next, we will further analyze the global modal perception mechanism of the Granular Ball, as well as the prior sources of threshold settings in the GBFF method.
V-A1 Granular Ball radius and significant pixel extraction
The radius setting of the Granular Ball affects the maximum boundary and splitting of the Granular Ball, which in turn impacts the classification of $SP$ and $NSP$, as shown in Fig.9 and Fig.10. When the radius is set too small, the sliding speed of the Granular Ball is too slow, resulting in a high number of iterations and significant computational time overhead, with $r_{POS}$ showing no obvious changes. However, for images with multiple exposure modes, fluctuations occur when the Granular Ball radius is set greater than 10, which is detrimental to subsequent modal perception. Therefore, we choose a Granular Ball radius of 10, balancing iteration speed while maintaining classification effectiveness.
Figure 9: The relationship between the radius of the Granular BALL and the number of iterations.
Figure 10: The relationship between the radius and $r_{POS}$ in the significant extraction algorithm for Granular BALL.
V-A2 Modal Perception
In the method proposed in this article, a threshold of 0.95 is set as the extreme modal perception condition for $r_{POS}$ to adjust the information ratio of multi-exposure images. As shown in Figure x, with a Granular Ball radius of 10, we statistically analyzed $r_{POS}$ for 100 images of each modality. It can be observed that the $r_{POS}$ obtained after Granular Ball generation classification exhibits significant clustering for images of different modalities. This allows us to set a threshold, which serves as a prior for global modal perception of the Granular Ball. Therefore, the threshold of 0.95 we set can encompass the vast majority of multi-exposure images without affecting images of other modalities. Meanwhile, multi-exposure images that are not perceived in some multi-modal images are treated as noise images by the neural network. Previous research has confirmed that this can enhance the robustness of the neural network.As shown in Table VII, after adding modal awareness, the values have all increased, indicating that the quality of the fused images has been further enhanced. From Fig.11, it can be observed that after adding modal perception, the suppression of the exposed areas has significantly improved, and the texture information is clearer.
Figure 11: In the case where the radius of the Granular BALL is 10, perform $r_{POS}$ evaluation on 100 pairs of images with different modalities.
TABLE VII: Quantitative comparison between Non-modal Perception and GBFF.
MEFB Metrics
Method EN MI SD SF Qab
Non-modal Perception 6.654 0.548 44.550 0.015 0.022
GBFF 6.947 0.678 46.655 0.028 0.031
V-B Priori guided training
In this section, we will conduct a comparative analysis using the priors provided by traditional methods and the Granular Ball prior to explore the role of prior guidance in neural network training[30][31][21]. Next, we will perform an ablation study on the weight guidance and supervised image structure guidance in the GBFF method. Finally, we will investigate the relationship between the number of training iterations and overfitting in Few-shot scenarios through experiments.
V-B1 Granular Ball prior and traditional method prior
A priori can effectively reduce the network’s demand for sample size. We will use images obtained from traditional methods as a priori to add to our network structure to explore the impact of different a priori on neural network training, and conduct qualitative and quantitative analyses. In this section, we will additionally introduce four traditional methods to provide a priori. These are: (c) the curvelet fusion method using curvelet transform, (d) the DTCWT method using dual-tree complex wavelet transform, (e) the GTT fusion method using guided filtering, and (h) the LP fusion method combining Laplace pyramid decomposition and sparse matrices. Since traditional methods do not have adaptive weight settings, we will set $r_{POS}$ and $r_{bnd}$ to 0.5 to balance the loss function. In the experiments, the number of iterations and parameter settings will be the same as those of the GBFF method. The fusion results are shown in Fig.12, which displays the fusion images from traditional methods and those using the image structure from traditional methods as a priori. It can be observed that the network with traditional methods as a priori can also achieve good results under few-shot conditions and can repair some of the poorer performance areas of the traditional methods’ a priori. Compared to the GBFF method, traditional methods cannot adaptively adjust the relationship between the neural network and the provided a priori images. Although the expressive power of the a priori supervision images provided by the GBFF method is not as strong as that of traditional methods in multi-focus fusion and medical image fusion, the final fusion images obtained after processing through the neural network are superior to those from traditional methods. The quantitative analysis of indicators is shown in Table VIII. Compared to traditional methods, the Few-shot method with traditional methods as a priori shows an increase in indicators, which validates the effectiveness of the Few-shot guided by a priori in this paper. In comparison to the GBFF method, it only achieved suboptimal results in the EN indicator for multi-exposure fusion, while achieving optimal results in other indicators. Therefore, the design and performance of the GBFF method are more suitable for general image fusion under few-shot conditions.
Figure 12: Compared to traditional algorithms and Few-shot methods that use traditional algorithms to fuse images as priors, images (d), (f), (g), and (i) are the fused images obtained from the prior-trained network (c), (e), (f), and (h) respectively, while (j) is the image obtained through Granular Ball Significant Extraction.
V-B2 Structural priors and weight-guided ablation
TABLE VIII: Comparison of fusion performance with and without prior enhancement across four datasets (MED, MFF, VIF, MEF). Methods suffixed with _prior include prior information.
Method MIF MFF VIF MEF
EN MI SD PSNR Qab EN SD SF AG Qab MI SD PSNR VIF Qab EN MI SD SF Qab
Curvelet 5.496 2.023 75.933 13.019 0.348 7.431 55.412 0.038 12.461 0.065 2.185 41.972 18.041 0.407 0.490 6.891 0.162 53.707 0.018 0.009
Curvelet_prior 5.263 2.623 79.704 13.985 0.531 7.499 57.614 0.034 11.923 0.139 2.528 40.340 17.695 0.394 0.496 5.627 0.376 22.175 0.012 0.011
Dtcwt 5.455 1.988 74.314 14.629 0.333 7.429 55.450 0.038 12.513 0.061 2.537 40.647 18.265 0.395 0.490 6.948 0.182 55.877 0.018 0.009
Dtcwt_prior 5.249 2.642 82.094 14.750 0.497 7.409 52.507 0.037 11.026 0.140 2.639 41.354 18.275 0.399 0.470 6.463 0.348 18.950 0.011 0.012
GFF 5.063 3.022 78.313 12.750 0.609 7.454 55.470 0.038 12.466 0.140 2.558 38.010 17.686 0.381 0.495 6.916 0.374 45.992 0.020 0.021
GFF_prior 5.235 2.717 82.361 12.775 0.567 7.483 58.097 0.040 12.866 0.136 2.256 40.214 18.213 0.400 0.492 6.543 0.359 38.896 0.019 0.024
Lp 4.796 2.177 77.944 13.356 0.385 7.432 55.775 0.038 12.574 0.041 2.446 40.184 18.651 0.393 0.496 6.452 0.215 53.145 0.019 0.012
Lp_prior 5.157 2.646 77.921 14.602 0.521 7.459 55.167 0.033 11.735 0.141 2.682 41.422 18.205 0.391 0.481 6.343 0.360 18.811 0.012 0.012
Supervise 4.834 3.731 85.419 11.655 0.619 7.379 52.650 0.023 7.408 0.083 3.170 39.891 18.238 0.402 0.494 6.640 0.639 40.718 0.020 0.022
GBFF 5.594 3.042 85.484 20.104 0.632 7.515 60.045 0.042 13.159 0.142 3.183 42.044 18.436 0.403 0.497 6.947 0.678 46.655 0.023 0.028
TABLE IX: Quantitative comparison of different ablation settings on MED, MFF, VIF, and MEF datasets. Each block of five metrics corresponds to one dataset.
Method MIF MFF VIF MEF
EN MI SD PSNR Qab EN SD SF AG Qab MI SD PSNR VIF Qab EN MI SD SF Qab
Non-weight 5.569 3.009 83.072 11.727 0.613 7.384 57.227 0.028 9.347 0.134 3.115 40.586 17.877 0.417 0.466 6.703 0.595 33.430 0.012 0.018
Non-prior 5.385 2.835 79.106 15.078 0.555 7.201 57.948 0.031 9.977 0.135 2.017 38.302 16.467 0.385 0.444 6.871 0.643 34.625 0.013 0.024
Non-prior and weight 5.601 2.042 59.766 7.363 0.493 7.339 55.494 0.033 11.386 0.093 1.925 33.251 8.706 0.378 0.273 6.879 0.611 42.430 0.017 0.019
OURS 5.594 3.042 85.484 20.104 0.632 7.515 60.045 0.042 13.159 0.142 3.183 42.044 18.436 0.403 0.497 6.947 0.678 46.655 0.028 0.031
In this section, we will gradually ablate the image prior and weight guidance to analyze their roles in few-shot scenarios for neural networks. After processing with Granular Ball, we obtain supervised images as the image prior for the neural network. As well as the $r_{pos}$ and $r_{bnd}$ of the prior images to serve as the confidence of the prior images, which will guide the loss function during the training of the neural network. Specifically, there are three cases: (c) Non-weight: with Granular Ball prior, no weight guidance ($r_{pos}$ and $r_{bnd}$ are both 0.5); (d) Non-prior: no Granular Ball prior, using a 0.5 weighted fusion image as a substitute; (e) Non-prior and weight: only using the neural network to fit the weights. As shown in Fig.13, methods lacking priors are not sensitive to salient features, while those lacking adaptive weight fusion of image edge information perform poorly. The fused images obtained without priors and weights exhibit severe distortion, making it difficult for the neural network to fit the expected fused image.From Table IX, it can be observed that the GBFF method achieved optimal results on most metrics, with the combined effect of structural priors and weight guidance resulting in the best performance of the fused images.
Figure 13: The fused image obtained from the structural prior provided by the ablation supervision image and the weight adjustment, as well as the fused image obtained from training the network without using prior and weight adjustment.
TABLE X: Ablation study on MED, MFF, VIF, and MEF datasets. Each column group corresponds to one dataset with five evaluation metrics.
Method MIF MFF VIF MEF
EN MI SD PSNR Qab EN SD SF AG Qab MI SD PSNR VIF Qab EN MI SD SF Qab
supervise 4.834 3.731 85.419 11.655 0.619 7.379 52.650 0.023 7.408 0.083 3.170 39.891 18.238 0.402 0.494 6.640 0.639 40.718 0.020 0.022
Non_lap 5.613 3.133 84.753 11.292 0.628 7.412 59.334 0.034 10.864 0.139 3.184 40.977 18.340 0.402 0.489 6.618 0.569 42.821 0.021 0.027
Non_loss 5.473 3.042 83.677 11.924 0.611 7.383 52.783 0.026 7.136 0.073 3.136 40.060 18.159 0.391 0.469 6.677 0.646 39.002 0.018 0.025
Non_lap and loss 5.506 3.053 84.060 11.538 0.604 7.370 52.046 0.024 7.127 0.087 3.162 39.654 18.195 0.399 0.464 6.433 0.488 39.210 0.020 0.025
OURS 5.594 3.042 85.484 20.104 0.632 7.515 60.045 0.042 13.159 0.142 3.183 42.044 18.436 0.403 0.497 6.947 0.678 46.655 0.028 0.031
V-B3 Selection of Iteration Count
In Few-shot learning, due to the use of supervised images as a prior for spatial structure, and since supervised images do not possess the expressiveness of real fused images, excessive iterations of the neural network can lead to overfitting. Early stopping of iterations is beneficial for maintaining higher expressiveness in the neural network. We evaluate the relationship between the fused images and the number of iterations by measuring the average values of PSNR, EN, MI, and SD of the output images for each iteration, as well as the changes in the loss function. From Fig.14, it can be observed that after 100 iterations, the various metrics of the fused images tend to stabilize, and the loss function also stabilizes. To avoid overfitting, we set the number of iterations to 100.
Figure 14: As the number of iterations increases, the changes in various evaluation metrics and the loss function.
V-C Loss function ablation
In the GBFF method, we describe the expected fused image as: the structural information of the supervised image plus salient information with confidence, ultimately supplemented by the boundary information and high-frequency information of the original fused image to compensate for areas where the saliency extraction confidence is insufficient. In this section, we will perform an ablation study on the loss function to verify the rationality of the loss function settings. In the previous section, it was already verified that networks lacking the structural prior of the supervised image find it difficult to learn the correct fused image information; therefore, we will not perform an ablation on $\mathcal{L}_{\text{ssim}}$ in this section. We will conduct experiments divided into three parts: (a) Non-$\mathcal{L}_{\text{Sobel}}$, (b) Non-$\mathcal{L}_{\text{Sobel}}$, and (c) Non-$\mathcal{L}_{\text{Lap}}$ and $\mathcal{L}_{\text{Sobel}}$. From Table X, it can be observed that the fused images obtained by networks lacking energy response and boundary response are less expressive than the supervised images. However, after adding $\mathcal{L}_{\text{Sobel}}$, the boundary extraction capability is enhanced, and after adding $\mathcal{L}_{\text{Lap}}$, the expressiveness of high-frequency detail information is improved. These changes in metric values align with our expectations for the loss function settings. The loss function settings of the GBFF method ensure that all metrics of the image reach optimal levels.
VI CONCLUSION
In this paper, we propose a general image fusion method for few-shot training using priors. To address the lack of priors in neural networks, we introduce a Granular Ball to perform saliency extraction on images. We design a Granular Ball algorithm that possesses modality-aware capabilities and adaptive saliency extraction in the brightness subspace, enabling it to provide priors for neural networks and adaptively adjust the neural network’s loss function. We validate the competitiveness and effectiveness of the proposed method in few-shot environments through qualitative and quantitative analysis against state-of-the-art (SOTA) methods. Next, we analyze the settings of the Granular Ball radius and modality awareness through experiments. Additionally, we demonstrate that neural networks guided by priors can perform image fusion in few-shot environments and compare different prior settings to validate the effectiveness of the Granular Ball prior. Finally, we conduct ablation experiments to analyze the effectiveness of various modules and loss functions set in this method. Although this paper presents a paradigm for few-shot image fusion and achieves effective results in typical image fusion domains, the Granular Ball’s settings only extract priors from the brightness subspace, which leads to the absence of other spaces. This may pose challenges of insufficient priors when applied to more tasks. In the future, we plan to develop joint inference algorithms for more subspaces through Granular Ball settings to achieve more precise and effective information extraction, and to apply Granular Ball priors to few-shot training in fields such as super-resolution and image segmentation.