### v2

# Rethinking Few-Shot Image Fusion: Granular Ball Priors Enable General-Purpose Deep Fusion

Abstract
In image fusion tasks, the absence of real fused images as priors presents a fundamental challenge. Most deep learning-based fusion methods rely on large-scale paired datasets to extract global weighting features from raw images, thereby generating fused outputs that approximate real fused images. In contrast to previous studies, this paper explores few-shot training of neural networks under the condition of having prior knowledge. We propose a novel fusion framework named GBFF, and a Granular Ball Significant Extraction algorithm specifically designed for the few-shot prior setting. All pixel pairs involved in the fusion process are initially modeled as a Coarse-Grained Granular Ball. At the local level, Fine-Grained Granular Balls are used to slide through the brightness space to extract Non-Salient Pixel Pairs, and perform splitting operations to obtain Salient Pixel Pairs. Pixel-wise weights are then computed to generate a pseudo-supervised image. At the global level, pixel pairs with significant contributions to the fusion process are categorized into the Positive Region, while those whose contributions cannot be accurately determined are assigned to the Boundary Region. The Granular Ball performs modality-aware adaptation based on the proportion of the positive region, thereby adjusting the neural network’s loss function and enabling it to complement the information of the boundary region. Extensive experiments demonstrate the effectiveness of both the proposed algorithm and the underlying theory. Compared with state-of-the-art (SOTA) methods, our approach shows strong competitiveness in terms of both fusion time and image expressiveness. Our code is publicly available at:
Index Terms:
Granular Ball, Image fusion, Few-shot, Deep learning
I Introduction
Due to the inherent limitations of individual sensors, images captured by a single sensor often fail to comprehensively represent all information within the field of view [1]. The goal of image fusion is to integrate complementary information from multiple sensors, merging images of different modalities to produce a more informative and comprehensive representation. For instance, visible light images captured at night cannot reveal thermal sources in dark regions, whereas infrared imaging can effectively highlight such thermal information [2]. By leveraging the strengths of various imaging modalities, fused images offer enhanced content and quality, making them highly valuable in applications such as medical imaging, surveillance, target recognition, and image segmentation [3].
In past research, significant progress has been made in various image fusion domains, including multi-focus image fusion, multi-exposure image fusion, infrared and visible light image fusion, and medical image fusion. Existing fusion methods can be broadly categorized into traditional approaches and deep learning-based approaches. Traditional methods typically rely on image decomposition techniques to extract features at different scales and in different domains for fusion [4]. However, as the demand for practical, real-world applications grows, these algorithmic designs have become increasingly complex and less adaptable. As a result, deep learning-based methods have gradually become mainstream. These approaches can be further divided into CNN-based, GAN-based, Transformer-based, and hybrid methods that integrate traditional techniques. Although these methods have demonstrated impressive performance, a key limitation remains: the lack of real fused images as ground truth in cross-modality fusion tasks. Consequently, most methods rely on large-scale datasets and require multiple iterations to generate results that approximate real fused images, which is often time-consuming. This presents practical challenges, such as long deployment periods and delayed model updates, in real-world applications.
To address the aforementioned challenges, we introduce Granular Balls into the domain of image fusion and design a granular saliency extraction algorithm with both global and local perception. This algorithm provides priors for the structural information and weight assignment during the neural network training, thus enabling few-shot learning for image fusion. Previous studies have shown that in most cases, pixel pairs with similar luminance across source images tend to carry redundant or overlapping information, while those with significant luminance differences are more likely to contain complementary information and exhibit salient behavior in the fusion process.[5, 6] Based on this prior, we operate in the luminance space, where pixel pairs with similar brightness are considered less contributive to fusion. At the local level, we employ fine-grained adaptive granular balls to extract pixel pairs with non-salient behavior and compute their weights based on the scale of granular ball boundaries.
This results in a pixel-level weight map used to generate a pseudo-supervised image. At the global level, salient pixel pairs are defined as the positive region, while pixel pairs lacking prior saliency are categorized as the boundary domain, whose ultimate behavior is determined by the neural network. The proportion of positive region pixels reflects the confidence of the pseudo-supervised image and serves as a prior for setting the loss function weight. Additionally, granular balls perceive modality-specific characteristics of the fused image through positive region, enabling targeted weight adjustment. Together, the pseudo-supervised image and the adaptive weight setting jointly guide the network toward rapid convergence. In summary, our contributions are as follows:
  • We propose an adaptive Granular Ball-based saliency extraction algorithm for general image fusion. This method enables not only the local extraction of salient pixel pairs and computation of their fusion weights, but also the global perception of modality characteristics, providing an effective prior for adaptive fusion under diverse scenarios.
  • To the best of our knowledge, this work is the first to achieve few-shot training for a general image fusion model across multiple modalities. By introducing adaptive Granular Balls as prior guidance, the proposed framework successfully performs fast and effective fusion with only 10 training image pairs per modality, establishing a novel paradigm for few-shot learning in the image fusion domain.
  • A complementary learning framework is established between Granular Balls and neural networks, where priors guide salient region learning while the network compensates for uncertain regions, resulting in fast convergence and high-quality fusion within 100 training epochs.
  • Compared with existing methods, the proposed approach achieves faster and more efficient image fusion. Extensive experiments demonstrate the effectiveness and rationality of the algorithm. Quantitative and qualitative results on public datasets show that our method attains competitive performance against several state-of-the-art (SOTA) techniques.
The remainder of this paper is organized as follows. Section II provides a brief overview of representative image fusion techniques and introduces the fundamentals of granular computing. Section III presents the proposed framework in detail, including its structural design and algorithmic implementation. Section IV compares our method with several state-of-the-art approaches, offering both qualitative and quantitative analysis. Section V conducts additional experimental evaluations, including ablation studies to validate the contribution of each component. Section VI concludes the paper and discusses potential future directions.
II Related Work
In this section, we will introduce the methods and developments of image fusion, as well as the theory and applications of granular computing.
II-A Image Fusion
Image fusion requires methods that can extract features from different modality images and highlight the effective information from each modality after fusion, thereby fully representing the real scene [7]. In past research, researchers primarily used methods such as multi-scale transformations, subspace analysis, and sparse representation. As the application scenarios became increasingly complex, some researchers began to combine these methods to enhance information extraction capabilities[8]. Early researchers used Laplacian pyramids (LP) to decompose images across different frequency bands, and later utilized dual-tree complex wavelet transform (DWCWT) to extract effective information through different filtering for image fusion. However, wavelets have poor edge extraction capabilities, leading researchers to propose the curvelet transform method, which effectively captures edge information. To enhance the representation of structural information in fusion, guided filtering (GFF) [9] was also applied to image fusion. Additionally, Liu et al. combined sparse matrix representation with traditional transformation methods to achieve good results in extracting effective information from fused images[10].
However, in real-world applications, image processing algorithms often suffer from poor generalizability as usage demands increase. This leads to increasingly complex designs, posing significant challenges for practical deployment [8]. With the rise of deep learning, Li et al. were the first to introduce it into image fusion [11], making it a mainstream approach in the field. Encoder–decoder networks, such as DeepFuse [12], demonstrated strong detail extraction capabilities in infrared and visible light fusion. Building upon this, IFCNN [13] utilized pre-trained networks for general-purpose fusion, while U2Fusion [14] designed a loss function that implicitly guides cross-modality fusion using saliency weights. In contrast to CNNs, Generative Adversarial Networks (GANs) use a discriminator to evaluate the generated results. FusionGAN [15] first introduced GANs to image fusion, and Ma et al. further proposed DDcGAN [16] by incorporating multiple discriminators to improve accuracy. Transformers, known for their long-range dependency modeling [17], were applied to image fusion by Ma et al. through SwinFusion [18], which treats image patches as tokens. Hybrid approaches combining traditional methods with deep learning have also emerged. For instance, MMIF-INET integrates wavelet decomposition with invertible neural networks for medical image fusion [19].Moreover, MetaFusion [20] uses meta-feature interaction for IR-VIS fusion. ZMFF [21, 22] introduces deep priors for sample-free multi-focus fusion. Despite these advances, many deep models—especially those based on CNNs and Transformers—require large-scale datasets to train effectively [23, 24]. Despite the promising results achieved by deep learning-based methods, practical deployment remains challenging. Approaches based on CNNs and Transformers typically require large-scale datasets to train models and extract parameters from source images. Meanwhile, sample-free methods such as ZMFF are computationally expensive, with each fused image requiring considerable processing time. Moreover, deep learning methods that rely on traditional priors struggle to determine an adaptive fusion scale. These limitations collectively hinder the rapid deployment and few-shot application of image fusion techniques
II-B Granular Ball Computing
Rough sets are widely used to model uncertainty in data, while ball granules offer a granulation-based approach to simplify fine-grained data, typically represented by a centroid and a radius within a spatial Granular Ball model [25, 26]. Building upon these foundations, researchers proposed the GBNRS method, which integrates granulation with approximation theory to handle continuous data more effectively [27]. Subsequently, Granular Ball clusters were introduced into clustering tasks to model instance relationships, leading to the Ball k-means algorithm, which significantly reduces spatial complexity during clustering [28]. Granular Ball computing has demonstrated strong potential in data analysis and uncertainty modeling. However, in the field of image processing, due to the lack of a unified modeling framework, existing ball granule-based methods have struggled to effectively capture relationships between image regions or modalities.
To address the above challenges, we map the original fused images into a brightness subspace and propose a Granular Ball Significant Extraction algorithm, which models pixel pairs based on their saliency in the brightness space. By integrating the Granular Ball mechanism with neural networks, we provide both structural priors and weight guidance, forming a complementary relationship between the Granular Ball and the network. This synergy enables few-shot training for general-purpose image fusion. The following sections will detail the overall structure and algorithmic implementation of the proposed method.
Figure 1: The overall framework of GBFF, where the dashed parts in the diagram are algorithms and modules used only during the training phase, and the solid parts are those needed in both the testing and training phases.
III Proposed Methods
In this section, we will introduce the overall framework process according to the training phase and the testing phase, and then provide a detailed explanation of the principles of the Granular Ball swarm algorithm, the network structure, and the configuration of the loss function.
III-A Framework
As shown in the Fig.1, the structural diagram of our framework is fully presented, including the settings of each module in the network. It is important to note that the dashed parts of the structural diagram are modules used only during the training phase, while the solid parts are used in both the testing and training phases. Here, SP refers to significant pixel pairs, and NSP refers to non-significant pixel pairs. We have two input paths, A and B, which can be interchanged freely. All input images will be converted to the YCbCr color space. In the network structure section, the numbers on the input paths and output paths represent the number of feature map channels, and all convolution modules in the network use 3x3 convolution kernels.
III-B Granular Ball generation
The brightness space of the original fused image carries the majority of the informative content, where the luminance of each pixel is responsible for representing image information.Since pixels appear in pairs during the fusion process, we construct fine-grained Granular Balls in the brightness space to serve as local-scale representations.
Definition 1. We define a granular ball $\mathcal{G}(\mu,r)$ in the brightness space as:
$\mathcal{G}(\mu,r)$ =$=\left\{x\in\mathbb{R}\,\middle|\,|x-\mu|\leq r\right\},$ $\mathcal{G}(\mu,r)\subseteq[0,255]$ (1)
Given a pair of source images $A$ and $B$, a pixel at position $(x,y)$ falls within the granular ball if:
$\begin{array}[]{c}(x,y)\in\mathcal{P}(\mu,r)\\[5.0pt] \text{s.t.}\quad A_{x,y}\in\mathcal{G}(\mu,r)\lor B_{x,y}\in\mathcal{G}(\mu,r)\end{array}$ (2)
III-C Granular Ball Significant Extraction
At the fine-grained level, in order to characterize the information difference between pixel pairs in the brightness space, we define the pixel pairs within each Granular Ball accordingly.
Definition 2. Each pixel pair $(A_{x,y},B_{x,y})$ from the source images is categorized into :
NSP: Both pixels fall within the granular ball.
$\begin{array}[]{c}(A_{x,y},B_{x,y})\in\mathcal{NSP}\\[5.0pt] \text{s.t.}\quad A_{x,y}\in\mathcal{G}(\mu,r)\land B_{x,y}\in\mathcal{G}(\mu,r)\end{array}$ (3)
SP: Only one of the pixels lies within the granular ball.
$\begin{array}[]{c}(A_{x,y},B_{x,y})\in\mathcal{SP}\\[5.0pt] \text{s.t.}\quad A_{x,y}\in\mathcal{G}(\mu,r)\land B_{x,y}\notin\mathcal{G}(\mu,r)\\ \quad\lor A_{x,y}\notin\mathcal{G}(\mu,r)\land B_{x,y}\in\mathcal{G}(\mu,r)\end{array}$ (4)
During the sliding process of fine-grained Granular Balls, there is insufficient prior knowledge to directly infer salient pixel pairs. As a result, the algorithm tends to first extract non-salient pixel pairs with similar brightness values. Meanwhile, we define the purity of each Granular Ball to describe its local state during the extraction process. Additionally, we introduce the concepts of the positive domain and the boundary domain to represent global information at a coarse-grained level.
Definition 3. Given a granular ball $\mathcal{G}(\mu,r)$, we define the purity of the granular ball as the proportion of non-salient pairs (NSP) within it:
$\text{Purity}(\mathcal{G})=\frac{|\mathcal{NSP}|}{|\mathcal{NSP}|+|\mathcal{SP}|}$ (5)
where, $|\mathcal{NSP}|$ and $|\mathcal{SP}|$ denote the number of pixel pairs classified as non-salient and salient, respectively, under the current granular ball $\mathcal{G}(\mu,r)$.
Definition 4.The pixel pairs in images are interpreted as follows:
  • The set of salient pairs (SP) corresponds to the positive region:
    $\mathcal{POS}(A,B)=\left\{(x,y)\,\middle|\,(A_{x,y},B_{x,y})\in\mathcal{SP}\right\}$ (6)
  • The set of non-salient pairs (NSP) corresponds to the boundary region:
    $\mathcal{BND}(A,B)=\left\{(x,y)\,\middle|\,(A_{x,y},B_{x,y})\in\mathcal{NSP}\right\}$ (7)
III-D Granular Ball Significant Extraction
As illustrated in the Fig.2, $NSP$(non-salient pixel pairs) are extracted locally through the sliding of coarse Granular Balls, while $SP$(salient pixel pairs) are extracted during the splitting process. For non-salient pairs, a greedy strategy is adopted to estimate their weights roughly, whereas salient pixel pairs aim to capture as much informative content as possible. By combining the sliding and splitting of fine-grained Granular Balls, all pixel pairs can be classified and assigned weights. The complete process of pixel pair extraction and weight computation is detailed in Algorithm 1.
    Algorithm 1:Granular Ball Significant Extraction
0:  Input images $A,B$
0:  Granular ball radius $r$, initial mask $X=0$, initial ranges $E_{1},E_{2}$
0:  Final fusion mask $X$, $\mathcal{SP}$, $\mathcal{NSP}$
1:  /* Function AssignSP$(E_{\text{low}},E_{\text{high}})$/
2:  for all $(x,y)\in\mathcal{SP}$ and $X_{x,y}=0$ do
3:     if $A_{x,y},B_{x,y}\in[E_{\text{low}},E_{\text{high}}]$ then
4:        if $A_{x,y}>B_{x,y}$ then
5:           $X_{x,y}=0.95$
6:        else
7:           $X_{x,y}=0.05$
8:        end if
9:     end if
10:  end for
11:  /* Function AssignNSP$(E_{\text{low}},E_{\text{high}})$/
12:  for all $(x,y)\in\mathcal{NSP}$ and $X_{x,y}=0$ do
13:     if $A_{x,y},B_{x,y}\in[E_{\text{low}},E_{\text{high}}]$ then
14:        $D_{1}=|\max(A_{x,y},B_{x,y})-E_{\text{high}}|$
15:        $D_{2}=|\max(A_{x,y},B_{x,y})-E_{\text{low}}|$
16:        $\omega=\max(D_{1},D_{2})/((E_{\text{high}}-E_{\text{low}})+10^{-6})$
17:        $\omega=\min(\omega,0.95)$
18:        if $A_{x,y}\geq B_{x,y}$ then
19:           $X_{x,y}=\max(\omega,0.05)$
20:        else
21:           $X_{x,y}=1-\max(\omega,0.05)$
22:        end if
23:     end if
24:  end for
25:  /* Main Procedure */
26:  Initialize $\mathcal{SP}=\varnothing$, $\mathcal{NSP}=\varnothing$
27:  $E_{1}=255$, $E_{2}=E_{1}-2r$
28:  while there exist unassigned pixels in $X$ do
29:     if $(E_{1}-E_{2})>6\cdot r$ then
30:        $\text{step}=\lfloor(E_{1}-E_{2})/2\rfloor$
31:        $\text{rangeLow}=E_{1}-\text{step}$,  $\text{rangeHigh}=E_{1}$
[…]
33:        $E_{1}=\text{rangeLow}$
34:     end if
35:     while $E_{1}>E_{2}$ and no unprocessed pixel in $[E_{2},E_{1}]$ do
36:        $E_{1}=E_{1}-1$
37:        if $E_{1}\leq E_{2}$ then
38:           $E_{2}=\max(E_{2}-2r,0)$
39:           $E_{1}=E_{2}$
40:        end if
41:     end while
42:     Call AssignSP$(E_{2},E_{1})$
43:     Call AssignNSP$(E_{2},E_{1})$
44:     $E_{2}=\max(E_{2}-2r,0)$
45:     $\mathcal{SP}^{\prime},\mathcal{NSP}^{\prime}=\{\text{newly unprocessed pixels in }[E_{2},E_{1}]\}$
46:     $\text{Purity}=\dfrac{|\mathcal{NSP}^{\prime}|}{|\mathcal{SP}^{\prime}|+|\mathcal{NSP}^{\prime}|}$
47:     if $\text{Purity}=1$ or $E_{2}<\delta$ then
48:        break
49:     end if
50:  end while
51:  return  $X,\mathcal{SP},\mathcal{NSP}$
The second part of the algorithm performs modality perception, as illustrated in Algorithm 2. The Granular Ball Significant Extraction yields global $NSP$ (non-salient pixel pairs) and $SP$ (salient pixel pairs), which are used to determine the ratio between the positive domain and the boundary domain. This positive domain ratio serves as a cue for modality perception, enabling adaptive weight adjustment. The $r_{\text{$POS$}}$ will be further explained in the experimental section. In this paper, modality perception is primarily applied to multi-exposure image fusion. Finally, the pseudo-supervised image and loss function weight guidance for the neural network are obtained through this process.
Figure 2: The process illustrates how non-salient pixel pairs are extracted through the sliding of Granular Balls, while salient pixel pairs are identified during the splitting phase, highlighting the dynamic changes in pixel pair classification throughout the extraction procedure.
    Algorithm 2: Modal Perception
0:  Final mask $X$, Classification $NSP$, $SP$,Input images $A,B$
0:  Supervise image Y , Output $r_{\text{$BND$}}$,$r_{\text{$POS$}}$
1:  Count total number of pixels $N$ in $X$
2:  Compute $POS$ ratio: $r_{\text{$POS$}}=\frac{N_{\text{$SP$}}}{N}$
3:  $r_{\text{$BND$}}$=$1-r_{\text{$POS$}}$
4:  if $r_{\text{POS}}>0.95$ then
5:     Identify all $(x,y)$ from $POS$
6:     for all such pixels $(x,y)$ do
7:        $X_{x,y}=0.5$
8:        $r_{\text{$BND$}},r_{\text{$POS$}}$ = $0.5,0.5$
9:     end for
10:  end if
11:  $Supervise=X*A+(1-X)*B$
12:  return  $Y$, $r_{\text{$BND$}}$,$r_{\text{$POS$}}$
III-E Multi-source input fusion strategy
The neural network trained by our method has the ability to respond to salient information and edges, as shown in Fig.3. By sequentially inputting multi-source images into the fully trained neural network, the network can extract salient information, thereby generating an image fused with salient information at each fusion step, ultimately resulting in a fused image that contains information from multiple source images.
Figure 3: Mutiple inputs fusion process
III-F Loss Function
Through the Granular Ball Significant Extraction algorithm, we obtain a pseudo-supervised image that provides structural priors, as well as $r_{\text{$POS$}}$ and $r_{\text{$BND$}}$ , which are used to guide the configuration of the loss function. The loss function is designed based on these components. We define the ideal fused image as comprising three key types of information: (1) the spatial structural information derived from the pseudo-supervised image, (2) the boundary information, where POS determines salient boundary features and BND preserves uncertain boundary cues from the original images, and (3) the high-frequency details contained in the original fused image. Accordingly, the total loss is divided into three components:
Figure 4: To better visualize the significance extraction process of Granular Balls during image weight computation, the image is divided into small patches of size 128×128. In subfigure (e) Weight Mask, the gray areas indicate regions whose weights have been adaptively adjusted.Subfigure (d) Supervise Image shows the pseudo-supervised image generated based on the weight mask.
a) Structural Similarity Loss:
$\mathcal{L}_{\text{ssim}}=1-\text{SSIM}(O,S)$ (8)
$\text{SSIM}(O,S)=\frac{(2\mu_{O}\mu_{S}+C_{1})(2\sigma_{OS}+C_{2})}{(\mu_{O}^{2}+\mu_{S}^{2}+C_{1})(\sigma_{O}^{2}+\sigma_{S}^{2}+C_{2})}$ (9)
where $\mu_{O}$, $\mu_{S}$ are local means, $\sigma_{O}^{2}$, $\sigma_{S}^{2}$ are variances, and $\sigma_{OS}$ is the covariance. Constants $C_{1}$ and $C_{2}$ stabilize the division.
b) Sobel Gradient Loss:
To guide the fusion process with edge-aware gradient information, we use the Sobel operator to extract gradients:
$G_{x}=\begin{bmatrix}-1&0&+1\\ -2&0&+2\\ -1&0&+1\end{bmatrix},\quad G_{y}=\begin{bmatrix}-1&-2&-1\\ 0&0&0\\ +1&+2&+1\end{bmatrix}$ (10)
The gradient magnitude is computed as:
$\text{Sobel}(I)=\sqrt{G_{x}^{2}+G_{y}^{2}}$ (11)
Since the confidence of the pseudo-supervised image is determined by the proportion of the positive region, denoted as $r_{\text{POS}}$, the boundary loss function is defined as:
$G_{\text{target}}=r_{\text{$POS$}}\cdot\text{Sobel}(S)+r_{\text{$BND$}}\cdot\left(\text{Sobel}(A)+\text{Sobel}(B)\right)$ (12)
where the $A$ and $B$ are original fused images.The Sobel loss minimizes the L1 difference between the output gradient and this guided target.The second part of loss function is:
$\mathcal{L}_{\text{Sobel}}=\|\text{Sobel}(O)-\text{Sobel}(G_{\text{target}})|_{1}$ (13)
c) Laplacian Loss:
The Laplacian filter is applied to all input and target images to extract high-frequency energy information:
$\Delta=\begin{bmatrix}0&1&0\\ 1&-4&1\\ 0&1&0\end{bmatrix}$ (14)
Therefore, the third part of the loss function is:
$\mathcal{L}_{\text{Lap}}=\|\Delta T-\Delta X\|_{1}+\|\Delta T-\Delta Y\|_{1}\$ (15)
Finally, we obtain the loss function for the entire network:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{ssim}}+\mathcal{L}_{\text{Lap}}+\mathcal{L}_{\text{Sobel}}\$ (16)
IV Experimental Results And Analysis
IV-A Experimental Setup
For the task of fusing infrared and visible light images(VIF), we selected 5 pairs of images from the MSRS[2] and TNO datasets[29] as the training set. For medical image fusion(MIF), we selected 10 pairs of images from the Harvard (PET-MRI) dataset[30] as the training set. For multi-exposure image fusion(MEF), we selected 10 pairs of images from the MEFB dataset[31] as the training set. Finally, for the multi-focus image fusion task(MFF), we selected 10 pairs of images from the Lytro dataset[32] as the training set. In the subsequent steps, we will use the images from the datasets that were not used in training as the test set. To validate generalizability, we will add the RoadScene dataset as a test set for the infrared and visible light image fusion task, and the MFI-WHU dataset[33] as a test set for multi-focus fusion.
During the training phase, to improve the robustness of the neural network, all input images are cropped to a size of 128×128 with a stride of 64 and a batch size of 16. To augment the dataset, we apply random flipping and shuffling within each batch. The total number of training epochs is set to 100. All experiments are conducted on a system equipped with an RTX 3090 GPU and an Intel i5-13600KF CPU.
TABLE I: Quantitative comparison of various medical image fusion methods is conducted on the MSRS, RoadScene, and TNO datasets.
Method MSRS RoadScene TNO
MI SD PSNR VIF Qab MI SD PSNR VIF Qab MI SD PSNR VIF Qab
BBDFusion 2.210 28.435 19.745 0.342 0.380 2.681 44.405 14.508 0.254 0.370 2.497 44.377 15.475 0.224 0.265
DeFusion 3.092 34.908 22.363 0.489 0.505 3.110 35.274 16.975 0.228 0.372 2.746 30.508 14.581 0.238 0.348
IFCNN 2.836 35.331 22.900 0.502 0.583 3.000 38.350 16.987 0.252 0.407 2.505 33.188 15.561 0.326 0.471
SDNET 0.570 27.364 12.118 0.019 0.026 0.988 36.904 12.451 0.026 0.034 2.661 26.462 11.163 0.023 0.046
SwinFusion 2.262 26.058 17.311 0.303 0.284 3.190 43.828 15.837 0.257 0.374 2.492 39.243 14.908 0.268 0.469
U2Fusion 2.262 26.058 17.311 0.303 0.284 3.025 39.830 14.945 0.195 0.390 2.110 28.594 15.665 0.281 0.337
MetaFusion 1.799 39.189 21.481 0.530 0.488 2.240 51.148 15.539 0.274 0.347 1.844 40.321 16.513 0.423 0.352
GBFF 3.306 38.499 22.388 0.533 0.591 3.351 46.254 17.333 0.324 0.425 2.892 41.378 15.566 0.354 0.475
IV-B Multimodal image fusion
In this section, we will compare our method with the following state-of-the-art (SOTA) methods for general image fusion: (c) DDBFusion[34], (d) DeFusion[35], (e) IFCNN, (f) SDNET[11], (g) SwinFusion, and (h) U2Fusion. To demonstrate the effectiveness of our method, we will introduce a dedicated SOTA method for each type of modality image fusion, specifically: (i) MetaFusion for infrared and visible light fusion, (j) MMIF-INET for MRI and PET image fusion, (k) ZMFF for multi-focus image fusion, and (l) SAMT-MEF[36] for multi-exposure image fusion.
In the quantitative analysis section, we will use the following evaluation metrics: MI (to assess the amount of information contained in the fused image), SD (to measure the contrast of the fused image), PSNR (to evaluate the reconstruction quality of the fused image), VIF (to assess the visual fidelity of the fused image), Qab (to evaluate the structural preservation of the fused image), EN (to assess the information entropy of the fused image), SF (to evaluate the spatial frequency of the fused image), and AG (to assess the clarity of the fused image). The higher the values of these evaluation metrics, the better the quality of the fused image. For ease of observation, we will use bold black to mark the optimal indicators and underline the sub-optimal indicators.
IV-B1 Infrared and visible image fusion
As shown in Fig.5, compared to the methods being compared, our method demonstrates a stronger advantage in the detail texture of visible light images and infrared features, resulting in fused images with less noise and more distinct boundaries. For quantitative analysis, we used MI , SD , PSNR, VIF, and Qab. The comparison results with various methods are shown in Table I. Our method shows a strong advantage in the amount of information, fidelity, and structural retention of the fused image, and it also achieved second place in other evaluation metrics. Overall, the fused images produced by our method demonstrate a more competitive performance, capable of completing fusion tasks with fewer samples, showcasing good generalization ability.
Figure 5: Select the fusion results of 3 typical images from TNO, RoadScene, and MSRS datasets for qualitative comparison.
IV-B2 Medical image fusion
The comparison results are shown in Fig.6. The fused images obtained from GBFF demonstrate a stronger advantage in texture details and edge performance, preserving the features of the PET images while also showcasing the texture characteristics of the MRI images. As shown in Table II, the GBFF method exhibits optimal results in the quantitative analysis of EN , MI, and PSNR values, indicating an advantage in the information content and contrast of the fused images. Additionally, the suboptimal SD and Qab values also indicate that the fused images obtained by the GBFF method are correlated with the original fused images.
TABLE II: Quantitative comparison of various medical image fusion methods is conducted on the Harvard (PET-MRI) dataset.
Harvard (PET-MRI) Metrics
Method EN MI SD PSNR Qab
BBDFusion 5.209 3.018 78.647 16.691 0.475
DeFusion 5.183 2.922 77.806 19.374 0.575
IFCNN 5.041 2.890 82.178 19.933 0.629
SDFusion 5.491 2.835 67.487 16.079 0.541
SwinFusion 5.233 2.548 83.406 18.956 0.600
U2Fusion 5.555 3.040 70.196 15.722 0.336
MMIF-INET 5.024 3.031 91.864 19.305 0.672
GBFF 5.594 3.042 85.484 20.104 0.632
Figure 6: Select the fusion results of 3 typical images from the Harvard (PET-MRI) dataset for qualitative comparison.
IV-B3 Multi-focus image fusion
As shown in Fig.7, the GBFF method benefits from the significant extraction of granules, resulting in fused images with higher boundaries and contrast compared to other methods, leading to clearer representations. As indicated in Tables III and IV, quantitative analyses on two datasets show that the GBFF method achieves optimal values in EN, SD, and Qab for fused images, indicating that the fused images maintain high contrast and information content while preserving the original fusion image information. Additionally, in terms of SF and AG , it achieved the best and second-best results on the two datasets, respectively. Overall, the GBFF method demonstrates stronger advantages and competitiveness compared to other methods.
TABLE III: Quantitative comparison of various medical image fusion methods is conducted on the Lytro dataset.
Lytro Metrics
Method EN SD SF AG Qab
BBDFusion 7.535 57.153 0.023 7.556 0.189
DeFusion 7.463 54.356 0.022 6.724 0.198
IFCNN 7.523 57.096 0.036 11.276 0.235
SDNET 7.482 55.270 0.033 10.369 0.244
SwinFusion 7.524 56.712 0.032 9.842 0.243
U2Fusion 7.460 58.390 0.023 7.073 0.207
ZMFF 7.524 56.954 0.035 11.003 0.059
GBFF 7.611 61.989 0.037 11.526 0.245
TABLE IV: Quantitative comparison of various medical image fusion methods is conducted on the MFI-WHU dataset.
MFI-WHU Metrics
Method EN SD SF AG Qab
BBDFusion 7.312 52.557 0.033 10.464 0.035
DeFusion 7.223 49.534 0.026 8.320 0.033
IFCNN 7.326 52.714 0.050 14.896 0.037
SDnet 7.288 51.255 0.045 14.744 0.038
SwinFusion 7.289 51.589 0.040 12.803 0.037
U2Fusion 7.226 53.326 0.026 8.230 0.032
ZMFF 7.295 51.994 0.045 14.780 0.035
GBFF 7.419 58.100 0.046 14.792 0.039
Figure 7: Select the fusion results of 3 typical images from the MFI-WHU and Lytro datasets for qualitative comparison.
IV-B4 Multi-exposure image fusion
Compared to other fusion methods, as shown in Fig.8, the GBFF method for multi-exposure image fusion focuses more on the representation of texture information, avoiding texture distortion like that caused by the IFCNN method and color distortion like that caused by the SAMT-MEF method. Additionally, in high-exposure environments, the GBFF method handles light sources more smoothly, preserving details without causing halo noise. Quantitative analysis is shown in Table V, where the GBFF method achieves optimal results in EN and Qab, indicating that the fused images obtained by the GBFF method can capture more information during fusion. The suboptimal results in MI, SD, and SF suggest that the fused images have a high amount of information and contrast.
TABLE V: Quantitative comparison of various medical image fusion methods is conducted on the MEFB dataset.
MEFB Metrics
Method EN MI SD SF Qab
BBDFusion 6.730 0.722 58.406 0.027 0.028
DeFusion 6.463 0.592 38.457 0.017 0.023
IFCNN 6.888 0.465 46.404 0.041 0.027
SDNET 6.623 0.668 40.202 0.023 0.025
SwinFusion 6.212 0.703 33.387 0.012 0.017
U2Fusion 6.739 0.744 39.674 0.013 0.022
SAMT-MEF 6.849 0.792 43.505 0.024 0.038
GBFF 6.947 0.678 46.655 0.028 0.031
Figure 8: Select the fusion results of 3 typical images from the MEFB dataset for qualitative comparison.
IV-C Efficiency Analysis
In this section, we will compare the efficiency of all methods using images of size 224*224 as a benchmark, and we will use the network’s training parameters (params), floating-point operations (Flops), and the time consumed to fuse a pair of images as metrics. As shown in Table VI, due to the fewer residual layers in the GBFF network and its simplicity, although it did not achieve the best results in terms of params, it achieved the optimal value in Flops and also performed second best in fusion time.
TABLE VI: Comparison of parameter size, computational cost, and inference time among different image fusion methods.
Method Params (M) FLOPs (G) Time (s)
BBDFusion 3.673 184.931 0.953
DeFusion 7.874 71.551 0.426
IFCNN 0.083 39.932 0.024
SDNET 0.044 25.638 0.056
SwinFusion 0.973 292.537 1.467
U2Fusion 0.659 405.174 0.143
SAMT-MEF 1.176 38.450 0.174
MMIF-INET 0.749 3.020 0.110
ZMFF 4.670 230.380 242.790
MetaFusion 9.350 24.630 0.220
GBFF 0.261 2.211 0.022
V Theoretical analysis and ablation experiments
V-A Analysis of Granular Ball Significant Extraction Algorithm
In this section, we will analyze the impact of the radius of the Granular Ball on the extraction of significant pixels. Next, we will further analyze the global modal perception mechanism of the Granular Ball, as well as the prior sources of threshold settings in the GBFF method.
V-A1 Granular Ball radius and significant pixel extraction
The radius setting of the Granular Ball affects the maximum boundary and splitting of the Granular Ball, which in turn impacts the classification of $SP$ and $NSP$, as shown in Fig.9 and Fig.10. When the radius is set too small, the sliding speed of the Granular Ball is too slow, resulting in a high number of iterations and significant computational time overhead, with $r_{POS}$ showing no obvious changes. However, for images with multiple exposure modes, fluctuations occur when the Granular Ball radius is set greater than 10, which is detrimental to subsequent modal perception. Therefore, we choose a Granular Ball radius of 10, balancing iteration speed while maintaining classification effectiveness.
Figure 9: The relationship between the radius of the Granular BALL and the number of iterations.
Figure 10: The relationship between the radius and $r_{POS}$ in the significant extraction algorithm for Granular BALL.
V-A2 Modal Perception
In the method proposed in this article, a threshold of 0.95 is set as the extreme modal perception condition for $r_{POS}$ to adjust the information ratio of multi-exposure images. As shown in Figure x, with a Granular Ball radius of 10, we statistically analyzed $r_{POS}$ for 100 images of each modality. It can be observed that the $r_{POS}$ obtained after Granular Ball generation classification exhibits significant clustering for images of different modalities. This allows us to set a threshold, which serves as a prior for global modal perception of the Granular Ball. Therefore, the threshold of 0.95 we set can encompass the vast majority of multi-exposure images without affecting images of other modalities. Meanwhile, multi-exposure images that are not perceived in some multi-modal images are treated as noise images by the neural network. Previous research has confirmed that this can enhance the robustness of the neural network[37].As shown in Table VII, after adding modal awareness, the values have all increased, indicating that the quality of the fused images has been further enhanced. From Fig.11, it can be observed that after adding modal perception, the suppression of the exposed areas has significantly improved, and the texture information is clearer.
Figure 11: In the case where the radius of the Granular BALL is 10, perform $r_{POS}$ evaluation on 100 pairs of images with different modalities.
TABLE VII: Quantitative comparison between w/o Modal Perception and GBFF.
MEFB Metrics
Method EN MI SD SF Qab
w/o Modal Perception 6.654 0.548 44.550 0.015 0.022
GBFF 6.947 0.678 46.655 0.028 0.031
V-B Priori guided training
In this section, we conduct a comparative analysis between traditional prior-based methods and the proposed Granular Ball priors, in order to investigate the role of prior guidance in neural network training. We then perform ablation studies on two key components of the GBFF framework: weight guidance and structural guidance from the pseudo-supervised image. Finally, we explore the impact of the number of training epochs under the few-shot setting.
V-B1 Granular Ball prior and traditional method prior
In previous studies on few-shot training, prior knowledge has been shown to effectively reduce the reliance of neural networks on large amounts of training data[38, 24, 39, 23]. Based on this observation, we incorporate prior images generated by traditional methods into our network as alternative guidance, aiming to explore how different types of priors influence the training process. Both qualitative and quantitative analyses are conducted to evaluate the impact of prior selection on fusion performance. In this section, we further introduce four traditional fusion methods to serve as priors: (c) the Curvelet fusion method based on curvelet transform, (d) the DTCWT method utilizing the dual-tree complex wavelet transform, (e) the GFF fusion method using guided filtering, and (h) the LP fusion method that combines Laplacian pyramid decomposition with sparse representation. Since traditional methods lack adaptive weight mechanisms, we manually set the values of $r_{\text{POS}}$ and $r_{\text{BND}}$ to 0.5 to balance the loss function. All other training settings, including the number of iterations and network parameters, remain consistent with those used in the GBFF method. The fusion results are illustrated in Figure 12, which compares the outputs of traditional methods and networks that use structure maps from traditional methods as priors. It can be observed that, even under few-shot conditions, networks guided by traditional priors still achieve promising results, and can partially compensate for the weaker regions present in the original traditional fusion outputs. However, compared to the GBFF method, traditional priors do not enable adaptive alignment between the network and the prior information. Although the structural priors extracted by traditional methods exhibit stronger feature representation in multi-focus and medical image fusion tasks, the final fused images produced by the GBFF-guided network outperform those obtained using traditional priors. A quantitative comparison is provided in Table VIII. Compared with standalone traditional methods, few-shot training with traditional priors achieves significant performance gains, verifying the effectiveness of the prior-guided few-shot training strategy proposed in this paper. When compared to the GBFF method, the traditional-prior variant only achieves slightly better results in the EN indicator for multi-exposure fusion, but the GBFF still consistently delivers the best performance across other metrics. This demonstrates that the design and performance of GBFF are better suited for general-purpose few-shot image fusion.
Figure 12: Compared to traditional algorithms and few-shot methods that use fused images from traditional algorithms as priors, images (d), (f), (g), and (i) represent the fused results generated by the networks trained with the corresponding priors from (c), (e), (f), and (h), respectively. Image (j) is the fused result generated based on the Granular Ball Significant Extraction.
V-B2 Structural priors and weight-guided ablation
TABLE VIII: Quantitative analysis is conducted on traditional methods and fusion results using traditional methods as priors.For each modality, the fusion performance metrics are averaged across all datasets.
Method MIF MFF VIF MEF
EN MI SD PSNR Qab EN SD SF AG Qab MI SD PSNR VIF Qab EN MI SD SF Qab
Curvelet 5.496 2.023 75.933 13.019 0.348 7.431 55.412 0.038 12.461 0.065 2.185 41.972 18.041 0.407 0.490 6.891 0.162 53.707 0.018 0.009
Curvelet_prior 5.263 2.623 79.704 13.985 0.531 7.499 57.614 0.034 11.923 0.139 2.528 40.340 17.695 0.394 0.496 5.627 0.376 22.175 0.012 0.011
Dtcwt 5.455 1.988 74.314 14.629 0.333 7.429 55.450 0.038 12.513 0.061 2.537 40.647 18.265 0.395 0.490 6.948 0.182 55.877 0.018 0.009
Dtcwt_prior 5.249 2.642 82.094 14.750 0.497 7.409 52.507 0.037 11.026 0.140 2.639 41.354 18.275 0.399 0.470 6.463 0.348 18.950 0.011 0.012
GFF 5.063 3.022 78.313 12.750 0.609 7.454 55.470 0.038 12.466 0.140 2.558 38.010 17.686 0.381 0.495 6.916 0.374 45.992 0.020 0.021
GFF_prior 5.235 2.717 82.361 12.775 0.567 7.483 58.097 0.040 12.866 0.136 2.256 40.214 18.213 0.400 0.492 6.543 0.359 38.896 0.019 0.024
Lp 4.796 2.177 77.944 13.356 0.385 7.432 55.775 0.038 12.574 0.041 2.446 40.184 18.651 0.393 0.496 6.452 0.215 53.145 0.019 0.012
Lp_prior 5.157 2.646 77.921 14.602 0.521 7.459 55.167 0.033 11.735 0.141 2.682 41.422 18.205 0.391 0.481 6.343 0.360 18.811 0.012 0.012
Supervise 4.834 3.231 85.419 11.655 0.619 7.379 52.650 0.023 7.408 0.083 3.170 39.891 18.238 0.402 0.494 6.640 0.639 40.718 0.020 0.022
GBFF 5.594 3.042 85.484 20.104 0.632 7.515 60.045 0.042 13.159 0.142 3.183 42.044 18.436 0.403 0.497 6.947 0.678 46.655 0.023 0.028
TABLE IX: A quantitative analysis is performed on the fusion results of ablation experiments involving prior guidance and weight guidance.For each modality, the fusion performance metrics are averaged across all datasets.
Method MIF MFF VIF MEF
EN MI SD PSNR Qab EN SD SF AG Qab MI SD PSNR VIF Qab EN MI SD SF Qab
w/o weight 5.569 3.009 83.072 11.727 0.613 7.384 57.227 0.028 9.347 0.134 3.115 40.586 17.877 0.417 0.466 6.703 0.595 33.430 0.012 0.018
w/o prior 5.385 2.835 79.106 15.078 0.555 7.201 57.948 0.031 9.977 0.135 2.017 38.302 16.467 0.385 0.444 6.871 0.643 34.625 0.013 0.024
w/o prior and weight 5.601 2.042 59.766 7.363 0.493 7.339 55.494 0.033 11.386 0.093 1.925 33.251 8.706 0.378 0.273 6.879 0.611 42.430 0.017 0.019
GBFF 5.594 3.042 85.484 20.104 0.632 7.515 60.045 0.042 13.159 0.142 3.183 42.044 18.436 0.403 0.497 6.947 0.678 46.655 0.028 0.031
In this section, we perform a step-by-step ablation of image priors and weight guidance to analyze their respective roles in few-shot neural network training. After processing with the Granular Ball, we obtain pseudo-supervised images as image priors, along with the corresponding $r_{\text{pos}}$ and $r_{\text{bnd}}$ values that reflect the confidence of these priors. These values are used to guide the loss function during training. Specifically, we consider the following three settings: (c) w/o Weight: Granular Ball prior is used, but no weight guidance is provided (i.e., $r_{\text{POS}}=r_{\text{BND}}=0.5$); (d) w/o Prior: No Granular Ball prior is used; instead, a simple 0.5-weighted fusion of the input images is used as a substitute prior; (e) w/o Prior and Weight: The network directly learns fusion without any prior or weight guidance. As shown in Figure 13, methods without prior information fail to capture salient features, while those without adaptive weight fusion struggle to preserve edge information. The fused results obtained without both priors and weights exhibit severe structural distortion, making it difficult for the network to fit the expected fused image. From Table IX, we observe that the full GBFF method achieves the best results across most evaluation metrics. This demonstrates that the combination of structural priors and weight guidance leads to the most effective fusion results.
Figure 13: The fused image obtained from the structural prior provided by the ablation supervision image and the weight adjustment, as well as the fused image obtained from training the network without using prior and weight adjustment.
TABLE X: quantitative analysis is conducted on the fused images obtained after the ablation of specific loss function components.For each modality, the fusion performance metrics are averaged across all datasets.
Method MIF MFF VIF MEF
EN MI SD PSNR Qab EN SD SF AG Qab MI SD PSNR VIF Qab EN MI SD SF Qab
supervise 4.834 3.231 85.419 11.655 0.619 7.379 52.650 0.023 7.408 0.083 3.170 39.891 18.238 0.402 0.494 6.640 0.639 40.718 0.020 0.022
w/o $\mathcal{L}_{\text{lap}}$ 5.613 3.033 84.753 11.292 0.628 7.412 59.334 0.034 10.864 0.139 3.184 40.977 18.340 0.402 0.489 6.618 0.569 42.821 0.021 0.027
w/o $\mathcal{L}_{\text{Sobel}}$ 5.473 3.042 83.677 11.924 0.611 7.383 52.783 0.026 7.136 0.073 3.136 40.060 18.159 0.391 0.469 6.677 0.646 39.002 0.018 0.025
w/o $\mathcal{L}_{\text{lap}}$ and $\mathcal{L}_{\text{Sobel}}$ 5.506 3.053 84.060 11.538 0.604 7.370 52.046 0.024 7.127 0.087 3.162 39.654 18.195 0.399 0.464 6.433 0.488 39.210 0.020 0.025
GBFF 5.594 3.042 85.484 20.104 0.632 7.515 60.045 0.042 13.159 0.142 3.183 42.044 18.436 0.403 0.497 6.947 0.678 46.655 0.028 0.031
V-B3 Selection of Train Epoch
In Few-shot learning, due to the use of supervised images as a prior for spatial structure, and since supervised images do not possess the expressiveness of real fused images, excessive iterations of the neural network can lead to overfitting. Early stopping of iterations is beneficial for maintaining higher expressiveness in the neural network. We evaluate the relationship between the fused images and the number of iterations by measuring the average values of PSNR, EN, MI, and SD of the output images for each iteration, as well as the changes in the loss function. From Fig.14, it can be observed that after 100 iterations, the various metrics of the fused images tend to stabilize, and the loss function also stabilizes.
Figure 14: As the number of iterations increases, the changes in various evaluation metrics and the loss function.
V-C Loss function ablation
In the GBFF method, the expected fused image is defined as a combination of the structural information from the pseudo-supervised image and salient information with confidence, further supplemented by the boundary information and high-frequency details from the original fused image to compensate for regions with low saliency confidence. In this section, we conduct an ablation study on the loss function to verify the rationality of its design. As confirmed in the previous section, when the network lacks structural priors from the pseudo-supervised image, it struggles to learn the correct fused image content. Therefore, we do not ablate the $\mathcal{L}_{\text{SSIM}}$ component in this part. The experiments are divided into three settings: (a) w/o $\mathcal{L}_{\text{Sobel}}$ , (b) w/o $\mathcal{L}_{\text{Lap}}$ , and (c) w/o both $\mathcal{L}_{\text{Sobel}}$ and $\mathcal{L}_{\text{Lap}}$. As shown in Table X, the fused images generated without boundary and detail supervision are significantly less expressive compared to those trained with the complete loss. Adding $\mathcal{L}_{\text{Sobel}}$ improves boundary sharpness, while including $\mathcal{L}_{\text{Lap}}$ enhances the network’s ability to reconstruct high-frequency textures. The corresponding changes in evaluation metrics confirm the effectiveness and necessity of each component. Overall, the complete loss function setting of GBFF ensures that the fused image achieves balanced performance across all metrics, and contributes to the superior expressiveness of the fusion results.
VI CONCLUSION
In this paper, we proposed a general-purpose image fusion framework that leverages prior information to enable few-shot training. To address the challenge of insufficient priors for neural networks, we introduced a Granular Ball-based saliency extraction algorithm, which operates in the brightness subspace with modality-aware and adaptive saliency extraction capabilities. This algorithm provides structural priors to the network and adaptively adjusts the loss function based on the extracted information. We conducted both qualitative and quantitative comparisons with state-of-the-art (SOTA) methods and demonstrated that our approach remains effective and competitive even under few-shot conditions. Additionally, we performed experimental analyses on the radius settings of the Granular Ball and its modality perception ability. The results show that prior-guided neural networks can achieve effective fusion in few-shot scenarios, and comparisons with different prior settings validate the effectiveness of the proposed Granular Ball prior. Finally, we conducted ablation studies to evaluate the contribution of each module and loss component in the proposed method. Although this work presents a novel paradigm for few-shot image fusion and achieves promising results across multiple representative fusion tasks, the current implementation extracts priors only from the brightness subspace, which may limit its effectiveness in more complex scenarios where information from other domains is needed. In future work, we plan to extend the Granular Ball mechanism to support joint inference across multiple subspaces, enabling more accurate and comprehensive information extraction. Furthermore, we aim to explore the use of Granular Ball priors in other few-shot tasks, such as super-resolution and image segmentation.